Let f(x)=−3x. The graph of f(x)​is transformed into the graph of g(x) by a vertical stretch of 4 units and a translation of 4 units right.
What is the equation for g(x)?

Answer by ikleyn(53763)

.
Let f(x)=−3x. The graph of f(x)​is transformed into the graph of g(x) by a vertical stretch
of 4 units and a translation of 4 units right. What is the equation for g(x)?
~~~~~~~~~~~~~

The new function g(x) is


    g(x) = 4*f(x-4) = 4*(-3(x-4)) = -12*(x-4).      ANSWER


I multiply f(x) by 4 to provide vertical stretch of 4 units, and I replace the argument "x" of the function f(x)

by (x-4) to translate the plot 4 units right.

Solved and explained.


===================

comment from student : Thank you! One question. Since it's translating to
the right, wouldn't it be -12 (x+4) though ?


My response :   NO.


It is a common  MISTAKE  to think that moving plot  4  units right requires using  (x+4)
instead of  x.   With such level of knowledge the  LOW  SCORE  is  GUARANTEED.


The correct rule is  EXACTLY  as it is stated in my post.


////////////////


To see many other similar and different problems solved,  look into the lesson
    - HOW TO write functions for transformed plots
in this site.

Learn the subject from there . . .